Surya2002nd:

prove that 1/tan3x+tanx-1/cot3x+cotx=cot4x

Surya2002nd:

Answer

ratnakermehta20:

We will use the following identity : \[(*)\[\tan(\alpha+\beta)=(\tan \alpha+\tan \beta)/(1-\tan \alpha \tan \beta).\]...\] \[The L.H.S =1/(\tan 3x+\tan x)-1/(\cot 3x+\cot x),\] \[=1/(\tan 3x+\tan x)-1/\left\{ 1/\tan 3x +1/\tan x \right\},\] \[1/(\tan 3x+\tan x)-1/\left\{ (\tan 3x+\tan x)/(\tan 3x \tan x) \right\},\] \[=1/(\tan 3x+\tan x) - (\tan 3x \tan x)/(\tan 3x+ \tan x),\] \[=(1-\tan 3x \tan x)/(\tan 3x+\tan x),\] \[=1/[(\tan 3x+\tan x)/(1-\tan 3x \tan x)],\] \[=1/[\tan(3x+x)]...[ by (*)],\] \[=1/\tan 4x,\] \[=\cot 4x,\] \[=The R.H.S.\] \[Hence, the Proof.\]